Edinburgh University Department of Computer Science

Student Demonstrator Short Curriculum Vitae

Name: Tim HEAP

Date of Birth: 25 July 1963

Degrees: awarding institution, dates etc.

B.A. (Hons.) Mathematics  The University of Cambridge 1986
M.Sc.  Theoretical Computer Science The University of Edinburgh 1994

Current Degree Registration: Ph.D. : Computer Assisted Proof LFCS, Department of Computer Science The University of Edinburgh

Previous tutoring/demonstrating experience: Tutoring experience outwith the University of Edinburgh ranges from Supervision of final year undergraduates at the University of Cambridge to tuition for G.C.S.E. re-sits. My previous experience of teaching undergraduates runs from pure algebra through to remedial mathematics for engineers.

I have tutored more specialised courses including Graph theory, Set Theory, Logic, Discrete Mathematics, and some Computational and Numerical Methods amongst others, as well as the usual courses in Algebra and Calculus.

Other relevant experience: My work on an extensive variety of computing projects extends back more than ten years and includes detailed experience of, for example, systems management and development, user and product support, software maintenance, network administration, and training and supervision of staff, as well as the usual analysis and programming jobs.

Additionally, my responsibilities have included the development and application of computer based methods for a full range of real engineering problems, such as design, modelling, analysis, economic evaluation, and preparation and presentation of material in reports to clients.

These tasks have required proficiency with a wide variety of programming languages and a comprehensive selection of commercial applications.

During my M.Sc. course I attended a variety of extra modules in areas other than those which I offered for assessment.